1)Suppose Demand for Apples (in bushels) is given by Q = 90-P and Supply is given by Q = P. The market for apples is dominated by a single, monopolistic firm "NYC Apples". Suppose you could regulate the market for Apples and impose a price ceiling. What price would maximize social welfare (combined producer and consumer surplus)?

2)Suppose Demand for Apples (in bushels) is given by Q = 90-P and Supply is given by Q = P. The market for apples is dominated by a single, monopolistic firm "NYC Apples".  How much more will consumers pay for Apples with NYC Apples being a monopoly compared to if it were a perfectly competitive market?
Answer #1

Answer:-1)

assuming, the MC of monopoly is the supply curve
MC=Q
and
The inverse demand curve
Q=90-P
P=90-Q
MR=90-2Q ........... The MR curve is double sloped than the demand curve
The monopoly produces at MR=MC
equating both
90-2Q=Q
3Q=90
Q=30
P=90-30=$60
the firm will charge $60

Answer:-2)

The monopoly produces at MR=MC
equating both
90-2Q=Q
3Q=90
Q=30

The firm will produce 30 units